OpenStudy (anonymous):

Matthew earns $8.50/h working at a grocery store. He receives a raise of 15% per hour. Which expression represents how much money Matthew will earn per hour after the raise? A. 8.50 x 3/20 B. 8.50 x 17/20 C. 8.50 x 23/20 D. 8.50 x 37/20

OpenStudy (aripotta):

well we can immediately rule out a and b cuz the fraction needs to be more than 1

OpenStudy (aripotta):

and just find out which of those fractions equal 115/100 (115% or 15% raise)

OpenStudy (anonymous):

That's what I can't figure out haha

OpenStudy (aripotta):

use a calculator, it's easy that way lol. 115/100 = 1.15

OpenStudy (anonymous):

I'm bad at precents, word problems, and fractions so I'm terrible at this.

OpenStudy (aripotta):

37/20 = 1.85 so that's not it

OpenStudy (aripotta):

23/20 = 1.15 so that's our answer, c.

OpenStudy (anonymous):

Thank you so much, omg. I have 3 more, can you help with those too? By the way, I love that little Hermoine monster avatar haha

OpenStudy (aripotta):

aw thank you :3 i made it myself. and yes, i'll help

OpenStudy (anonymous):

Seriously? That awesome. And alright let me post the question :)

OpenStudy (aripotta):

are you understanding this better tho?